You are graduating from college at the end of this semester and after reading the The Business of Life box in this​ chapter, you have decided to invest $4,700 at the end of each year into a Roth IRA for the next 42 years. If you earn 6 percent compounded annually on your​ investment, how much will you have when you retire in 42 years? How much will you have if you wait 10 years before beginning to save and only make 32 payments into your retirement​ account? ​(Round to the nearest​ cent.)
Answer #1

Future value=Annuity/rate*((1+rate)^t-1)

1.
=4700/6%*(1.06^42-1)=826967.559475451

2.
=4700/6%*(1.06^32-1)=427181.956747685